CHEMICAL AND SPECTROSCOPIC APPLICATIONS OF ULTRACOLD IONS IN ION TRAPS

\underline{STEFAN WILLITSCH}, Department of Chemistry, University College London, 20 Gordon Street, London WC1H 0AJ, United Kingdom.

Over the past years, ultracold (T < 10 mK) atomic and molecular ions prepared by laser and sympathetic cooling in ion traps have found an increasing number of applications in molecular physics^{a}, chemistry^{b} and spectroscopy^{c}. At low temperatures the ions undergo a phase transition from a ”gaseous” state to an ordered structure in the trap. Owing to their strong localisation single ions can be observed, addressed and manipulated opening up perspectives for new gas-phase experiments with single-particle sensitivity. In this talk, novel experimental schemes for reactive-scattering studies and high-resolution spectroscopy based on ultracold ions will be discussed. A new method to investigate reactive collisions between laser-cooled ions and velocity-selected neutral molecules will be introduced. The technique is particularly suited to measure ion-molecule reaction rates as a function of temperature down to ≈ 1 K. First results on the Ca^{+} + CH_{3}F → CaF^{+} + CH_{3} reaction will be presented. In the second part of the talk, prospects for the state-selected generation of cold molecular ions and their application in high-resolution spectroscopic experiments will be explored.

aM. Drewsen, I. Jensen, J. Lindballe, N. Nissen, R. Martinussen, A. Mortensen, P. Staanum and D. Voigt, Int. J. Mass Spectrom.

229 (2003), 203

bS. Willitsch, M. T. Bell, A. D. Gingell, S. R. Procter and T. P. Softley, Phys. Rev. Lett. 100 (2008), 043203

cP. O. Schmidt, T. Rosenband, C. Langer, W. M. Itano, J. C. Bergquist and D. J. Wineland, Science 309 (2005), 749